How many moles are in 3.01 x 1024 formula units of RbNO3?
1
Expert's answer
2020-03-04T07:59:39-0500

One mole of RbNO3 has 6.022 "\times" 1023 molecules or units

6.022"\times" 1023 molecule of RbNO3 is equal to 1 mole

so, 3.01 "\times" 1024 molecules is equal to "=\frac{3.01 \times 10^{24}}{6.022\times10^{23}}= 4.998 mole"

which is approx equal to 5 mole

so total moles will be 5 mole